# Soliton solutions to the Sawada–Kotera equation

### v1

@@ 1 Introduction
In this paper we are interested in soliton solutions to an integrable system of coupled fifth-order nonlinear partial differential equations and, in particular, one of its special cases, namely, the Sawada–Kotera equation [26]. We construct such solutions by solving the inverse scattering problem for a third-order differential equation in the reflectionless case. The solution to the inverse scattering problem is obtained by solving a related Riemann–Hilbert problem using input consisting of the bound-state poles of a transmission coefficient and the bound-state dependency constants. Using the time-evolved dependency constants in the input data set, we show that the solution to the inverse problem yields soliton solutions to the aforementioned integrable nonlinear system and also soliton solutions to the Sawada–Kotera equation. This method explains the physical origin of the $2\mathbf{N}$ real parameters appearing in the relevant $\mathbf{N}$-soliton solution formula obtained by the bilinear method [17] of Hirota, by relating $\mathbf{N}$ of the real parameters to the bound-state energy values and relating the remaining $\mathbf{N}$ real parameters to the bound-state dependency constants.
@@
In the inverse scattering transform method, a time-evolved potential is constructed from the time-evolved scattering data. A soliton solution corresponds to a time-evolved potential corresponding to a reflectionless scattering data set. A reflectionless scattering data set comprises the bound-state information only. The bound-state information in turn consists of the bound-state energies and the bound-state normalization constants, or equivalently it consists of the bound-state poles of a transmission coefficient and a bound-state dependency constant for each bound-state pole. For example, an $\mathbf{N}$-soliton solution to the Sawada–Kotera equation is a real-valued function of $x$ and $t$ and it contains $2\mathbf{N}$ real parameters. As our method indicates, $\mathbf{N}$ of those real parameters identify the location of the bound-state poles of a transmission coefficient and the remaining $\mathbf{N}$ real parameters each identify the bound-state dependency constants associated with the bound-state poles.
Our paper is organized as follows. In Section 2 we introduce the integrable system (2.6) of coupled nonlinear partial differential equations associated with the third-order linear equation (2.8). We indicate how the Sawada–Kotera equation arises as a special case of (2.6) by uncoupling the integrable system (2.6). In Section 3 we provide a summary of the direct scattering problem for (2.8) in the reflectionless case. This is done by introducing the relevant solutions to (2.8) at each $k$-value in the complex $k$-plane. We indicate how the left and right transmission coefficients are related to the spacial asymptotics of two relevant solution to (2.8). In Section 4 we introduce the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k),$ and we describe the bound-state dependency constant for each bound-state pole. For simplicity we assume that each bound-state pole is simple. In Section 5 we describe the solution to the inverse scattering problem for (2.8) in the reflectionless case. This is done by providing some explicit expressions for the potentials $Q$ and $P$ in terms of the input data set consisting of $T_{\text{\rm{l}}}(k),$ the bound-state poles of $T_{\text{\rm{l}}}(k),$ and the bound-state dependency constants specified for the bound-state poles. In the case of the recovery of the potentials $Q$ and $P$ that depend on the parameter $t,$ the use of time-evolved dependency constants yields soliton solutions to the corresponding integrable system. In Section 6, by imposing the appropriate restrictions on a certain solution to the nonlinear system (2.6), we obtain the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). Finally, in Section 7, we illustrate the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation when $\mathbf{N}=1,2,3.$
@@ 2 The integrable system and the Sawada–Kotera equation
where the subscripts denote the corresponding partial derivatives. It is used as a model to describe the propagation of surface water waves in long, narrow, shallow canals, similar to the modeling used by the KdV (Korteweg–de Vries) equation [19]. For both the KdV equation and the Sawada–Kotera equation, we assume that the solutions are real valued. Hence, the quantity $Q$ appearing in (2.1) is assumed to be real valued. Compared to third-order nonlinear KdV equation, the Sawada–Kotera equation takes into account higher-order nonlinear and dispersive effects. In Section 7 of our paper, we remark on some different behaviors of soliton solutions to the Sawada–Kotera equation and to the KdV equation.
The Sawada–Kotera equation is integrable in the sense of the inverse scattering transform. This is due to the fact that it has a Lax pair, i.e. it is related to two linear differential operators $L$ and $A$ satisfying the Lax operator equation [20]
@@
where $L_{t}$ is obtained by applying the partial $t$-derivative on $L.$ By using the corresponding Lax pair $(L,A)$ on the left-hand side of (2.2), that left-hand side becomes the zero multiplication operator when (2.1) holds. There are actually two different Lax pairs for the Sawada–Kotera equation (2.1). The first pair $(L,A)$ is given by
@@
where the potentials $Q$ and $P$ are independent of each other and are allowed to take complex values. The Lax pair in (2.3) is obtained from (2.5) when $Q$ is real valued and we have $P\equiv 0,$ and the Lax pair in (2.4) is obtained when $Q$ is real valued and we have $P=Q_{x}$. The general Lax pair $(L,A)$ given in (2.5) satisfies the Lax operator equation (2.2) provided the potentials $Q$ and $P$ appearing in (2.5) satisfy the fifth-order coupled nonlinear partial differential equations
@@
In order to uncouple the nonlinear system (2.6), from the last three terms of the first coupled equation in (2.6) we see that we must have
@@
which shows that the uncoupling in (2.6) occurs when $P\equiv 0$ or when $P=Q_{x}.$ For each of those two cases, one can directly verify that the left-hand side of the second coupled equation in (2.6) vanishes identically. Hence, we confirm that the use of $P\equiv 0$ or $P=Q_{x}$ in (2.6) yields (2.1).
@@
$\psi^{\prime\prime\prime}+Q(x)\,\psi^{\prime}+P(x)\,\psi=k^{3}\,\psi,\qquad x\in\mathbb{R}.$ (2.8)
We suppress the dependence on $t$ for the potentials $Q$ and $P,$ and we write $Q(x)$ and $P(x),$ respectively, instead of $Q(x,t)$ and $P(x,t).$ The adjoint equation [3] for (2.8) is given by
@@
$\overline{Q}(x):=Q(x)^{\ast},\qquad\overline{P}(x):=Q^{\prime}(x)^{\ast}-P(x)^{\ast},\qquad x\in\mathbb{R},$ (2.10)
where we use an asterisk for complex conjugation and use an overbar to denote the quantities associated with the adjoint equation. Comparing (2.8) and (2.9), with the help of (2.10) we observe that, when the potential $Q$ is real valued and the potential $P$ is zero, the adjoint potential $\overline{Q}$ is equal to $Q$ and the adjoint potential $\overline{P}$ becomes equal to $Q_{x}.$ In that case, the equation in (2.8) reduces to
@@ 3 The direct scattering problem in the reflectionless case
where $\arg[k]$ denotes the argument function taking values in the interval $(-2\pi/3,4\pi/3).$ We use $\overline{\Omega_{1}},$ $\overline{\Omega_{2}},$ $\overline{\Omega_{3}},$ $\overline{\Omega_{4}}$ to denote the closures of the open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4},$ respectively, where we recall that each closure is obtained by adding the boundaries to the corresponding open sector.
@@
$T_{\text{\rm{r}}}(k)=\displaystyle\frac{1}{T_{\text{\rm{l}}}(k)},\qquad k\in\mathbb{C}.$
It is known [3] that (2.8) has the particular solutions $m(k,x)$ and $n(k,x)$ with the respective $k$-domains $\overline{\Omega_{2}}$ and $\overline{\Omega_{3}},$ and in the reflectionless case they satisfy the spacial asymptotics given by
@@
Since $k$ appears as $k^{3}$ in (2.8), it follows that $\psi(zk,x)$ and $\psi(z^{2}k,x)$ are both solutions to (2.8) whenever $\psi(k,x)$ is a solution. The $k$-domain of $\psi(zk,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $2\pi/3$ radians around the origin of the complex $k$-plane. Similarly, the $k$-domain of $\psi(z^{2}k,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $4\pi/3$ radians. This allows us to determine three linearly independent solutions to (2.8) at each $k$-value in the complex $k$-plane. Toward that goal, we divide $\Omega_{1}$ into the open subsectors $\Omega_{1}^{\text{\rm{up}}}$ and $\Omega_{1}^{\text{\rm{down}}}$ by using the negative real axis $\mathbb{R}^{-},$ and we divide the open sector $\Omega_{3}$ into the open subsectors $\Omega_{3}^{\text{\rm{down}}}$ and $\Omega_{3}^{\text{\rm{up}}}$ by using the positive real axis $\mathbb{R}^{+}.$ On the middle plot of Figure 3.1, we show the partitioning of the complex $k$-plane into the six open sectors $\Omega_{1}^{\text{\rm{up}}},$ $\Omega_{1}^{\text{\rm{down}}},$ $\Omega_{2},$ $\Omega_{3}^{\text{\rm{down}}},$ $\Omega_{3}^{\text{\rm{up}}},$ and $\Omega_{4}.$ We use -$\mathbb{R}^{-}$ and $\mathbb{R}^{+},$ respectively, to denote the directed half lines parametrized as
@@
For each fixed real values of $x$ and $t,$ the large $k$-asymptotics of the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ are, respectively, given by [3]
@@
Using (3.10)–(3.13), we express the potentials $Q$ and $P$ in terms of $u_{1}(x),$ $u_{2}(x),$ $v_{1}(x),$ $v_{2}(x)$ as
@@ 4 The bound states and the bound-state dependency constants
A bound state corresponds to a nontrivial solution to (2.8) which is square integrable in $x\in\mathbb{R}.$ If a bound state occurs at $k=k_{j}$ in the complex $k$-plane, then the number of linearly independent square-integrable solutions to (2.8) at $k=k_{j}$ determines the multiplicity of that bound state. In our paper, we only consider simple bound states, where the multiplicity of each bound state is equal to $1.$ We refer the reader to [4, 5, 6] for the treatment of bound states with multiplicities for various differential or difference equations. The analysis of solutions to (2.8) at $k=0$ is more challenging. When we analyze bound states for (2.8), since we assume that the potentials $Q$ and $P$ belong to the Schwartz class in $x\in\mathbb{R},$ we exclude the possibility of a bound state at $k=0.$
The bound states occur at the zeros of $T_{\text{\rm{l}}}(k)^{-1}$ in $\Omega_{1}^{\text{\rm{down}}}$ or $\Omega_{1}^{\text{\rm{up}}}.$ Without loss of generality, it is enough to consider the bound states at the poles of $T_{\text{\rm{l}}}(k)$ in $\Omega_{1}^{\text{\rm{down}}}.$ It turns out [3] that the analysis of the bound state at $k=k_{j}$ with $k_{j}\in\Omega_{1}^{\text{\rm{down}}}$ yields useful information also about the bound state at $k=k_{j}^{\ast}$ with $k_{j}^{\ast}\in\Omega_{1}^{\text{\rm{up}}}.$ We recall that we use an asterisk to denote complex conjugation.
@@
where we have the determinant of the relevant $3\times 3$ matrix on the right-hand side. The $3$-Wronskian of any three solutions to (2.8) at any particular $k$-value is zero if and only if those three solutions are linearly dependent. Furthermore, because of the absence of $\psi^{\prime\prime}$ in (2.8), the $3$-Wronskian of those three solutions is independent of $x$ and its value can be evaluated at any particular $x$-value. For example, using their asymptotics as $x\to\pm\infty,$ we evaluate the $3$-Wronskian of $f(k,x),$ $g(zk,x),$ and $n(z^{2}k,x)$ in $\overline{\Omega_{1}^{\text{\rm{down}}}}$ as
@@
Let us assume that the nonzero complex constant $k_{j}$ is located in the open sector $\Omega_{1}^{\text{\rm{down}}}$ and that it corresponds to a bound state. Then, we have $T_{\text{\rm{l}}}(k_{j})^{-1}=0.$ From (4.1) we see that the three solutions $f(k_{j},x),$ $g(zk_{j},x),$ and $n(z^{2}k_{j},x)$ to (2.8) become linearly dependent. This allows us [3] to express $f(k_{j},x)$ as a linear combination of $g(zk_{j},x)$ and $n(z^{2}k_{j},x)$ as
@@
for some complex constants $D(k_{j})$ and $W(k_{j}).$
@@
where the nonzero complex constant $W(k_{j})$ corresponds to the dependency constant at the bound state with $k=k_{j}.$ On the other hand, from the analysis in Section 3 of [3] it follows that $W(k_{j})=0$ if we have $\arg[k_{j}]\in[7\pi/6,4\pi/3).$ In that case, from (4.2) with $W(k_{j})=0$ we get
@@
with the nonzero complex constant $D(k_{j})$ corresponding to the dependency constant at the bound state with $k=k_{j}.$ We remark that, if the potentials $Q$ and $P$ in (2.8) depend also on the parameter $t,$ then the solutions to (2.8) also depend on $t.$ Consequently, the dependency constants $W(k_{j})$ and $D(k_{j})$ appearing in (4.3) and (4.4) also depend on $t.$ For further analysis related to the bound states of (2.8), we refer the reader to [3].

### v2

@@ 1 Introduction
In this paper we are interested in soliton solutions to an integrable system of coupled fifth-order nonlinear partial differential equations and, in particular, one of its special cases, namely, the Sawada–Kotera equation [26]. We construct such solutions by solving the inverse scattering problem for a third-order linear differential equation in the reflectionless case. The solution to the inverse scattering problem is obtained by solving a related Riemann–Hilbert problem using input consisting of the bound-state poles of a transmission coefficient and the bound-state dependency constants. Using the time-evolved dependency constants in the input data set, we show that the solution to the inverse problem yields soliton solutions to the aforementioned integrable nonlinear system and also soliton solutions to the Sawada–Kotera equation. This method explains the physical origin of the $2\mathbf{N}$ real parameters appearing in the relevant $\mathbf{N}$-soliton solution formula obtained by the bilinear method [17] of Hirota, by relating $\mathbf{N}$ of the real parameters to the bound-state poles and relating the remaining $\mathbf{N}$ real parameters to the bound-state dependency constants.
@@
In the inverse scattering transform method, a time-evolved potential is constructed from the time-evolved scattering data. A soliton solution corresponds to a time-evolved potential corresponding to a reflectionless scattering data set. A reflectionless scattering data set comprises the bound-state information only. The bound-state information in turn consists of the bound-state poles of a transmission coefficient and a bound-state dependency constant for each bound-state pole. For example, an $\mathbf{N}$-soliton solution to the Sawada–Kotera equation is a real-valued function of $x$ and $t$ and it contains $2\mathbf{N}$ real parameters. As our method indicates, $\mathbf{N}$ of those real parameters identify the locations of the bound-state poles of a transmission coefficient and the remaining $\mathbf{N}$ real parameters identify the bound-state dependency constants associated with the bound-state poles.
Our paper is organized as follows. In Section 2 we introduce the fifth-order integrable system (2.6) of coupled nonlinear partial differential equations associated with the third-order linear equation (2.8). We indicate how the Sawada–Kotera equation arises as a special case of (2.6) by uncoupling the integrable system (2.6). In Section 3 we provide a summary of the direct scattering problem for (2.8) in the reflectionless case. This is done by introducing three relevant solutions to (2.8) at each $k$-value in the complex $k$-plane. We indicate how the left and right transmission coefficients are related to the spacial asymptotics of two of the relevant solutions to (2.8). In Section 4 we introduce the bound-state poles of the left transmission coefficient $T_{\text{\rm{l}}}(k),$ and we describe the bound-state dependency constant for each bound-state pole. For simplicity we assume that each bound-state pole is simple. In Section 5 we describe the solution to the inverse scattering problem for (2.8) in the reflectionless case. This is done by providing some explicit expressions for each of the potentials $Q$ and $P$ in terms of the input data set consisting of $T_{\text{\rm{l}}}(k),$ the bound-state poles of $T_{\text{\rm{l}}}(k),$ and the bound-state dependency constants. In the case of the recovery of the potentials $Q$ and $P$ that depend on the parameter $t,$ the use of time-evolved dependency constants yields soliton solutions to the corresponding integrable system. In Section 6, by imposing the appropriate restrictions on a certain solution to the nonlinear system (2.6), we obtain the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation (2.1). Finally, in Section 7, we illustrate the $\mathbf{N}$-soliton solution to the Sawada–Kotera equation when $\mathbf{N}$ takes the values of $1,$ $2,$ and $3.$
@@ 2 The integrable system and the Sawada–Kotera equation
where $\mathbb{R}$ is the real axis and the subscripts denote the corresponding partial derivatives. It is used as a model to describe the propagation of surface water waves in long, narrow, shallow canals, similar to the modeling used by the KdV (Korteweg–de Vries) equation [19]. For both the KdV equation and the Sawada–Kotera equation, we assume that the solutions are real valued. Hence, the quantity $Q$ appearing in (2.1) is assumed to be real valued. Compared to the third-order nonlinear KdV equation, the Sawada–Kotera equation takes into account higher-order nonlinear and dispersive effects. In Section 7 of our paper, we remark on some similar behaviors of soliton solutions to the Sawada–Kotera equation and to the KdV equation.
The Sawada–Kotera equation is integrable in the sense of the inverse scattering transform method. This is due to the fact that it has a Lax pair, i.e. it is related to two linear differential operators $L$ and $A$ satisfying the Lax operator equation [20]
@@
where $L_{t}$ is obtained by applying the partial $t$-derivative on $L.$ By using the corresponding Lax pair $(L,A)$ on the left-hand side of (2.2), that left-hand side becomes the zero multiplication operator when (2.1) holds. There are actually two distinct Lax pairs for the Sawada–Kotera equation (2.1). The first pair $(L,A)$ is given by
@@
where the potentials $Q$ and $P$ are independent of each other and are allowed to take complex values. The Lax pair in (2.3) is obtained from (2.5) when $Q$ is real valued and we have $P\equiv 0,$ and the Lax pair in (2.4) is obtained when $Q$ is real valued and we have $P=Q_{x}$. The general Lax pair $(L,A)$ given in (2.5) satisfies the Lax operator equation (2.2) provided the potentials $Q$ and $P$ appearing in (2.5) satisfy the coupled system of two fifth-order nonlinear partial differential equations
@@
In our paper we assume that $Q$ and $P$ belong to the Schwartz class in $x\in\mathbb{R}$ for each fixed $t.$ In order to uncouple the nonlinear system (2.6), from the last three terms on the left-hand side of the first coupled equation in (2.6) we see that we must have
@@
which shows that the uncoupling in (2.6) occurs when $P\equiv 0$ or when $P=Q_{x}.$ For each of those two cases, one can directly verify that the left-hand side of the second coupled equation in (2.6) vanishes identically when the first equation holds. Hence, we confirm that the use of $P\equiv 0$ or $P=Q_{x}$ in (2.6) yields (2.1).
@@
$\psi^{\prime\prime\prime}+Q(x)\,\psi^{\prime}+P(x)\,\psi=k^{3}\,\psi,\qquad x\in\mathbb{R},$ (2.8)
where the prime denotes the $x$-derivative. We suppress the dependence on $t$ for the potentials $Q$ and $P,$ and we write $Q(x)$ and $P(x),$ respectively, instead of $Q(x,t)$ and $P(x,t).$ The adjoint equation [3] for (2.8) is given by
@@
$\overline{Q}(x)=Q(x)^{\ast},\qquad\overline{P}(x)=Q^{\prime}(x)^{\ast}-P(x)^{\ast},\qquad x\in\mathbb{R},$ (2.10)
where we use an asterisk for complex conjugation and use an overbar to denote the quantities associated with the adjoint equation (2.9). Comparing (2.8) and (2.9), with the help of (2.10) we observe that, when the potential $Q$ is real valued and the potential $P$ is zero, the adjoint potential $\overline{Q}$ is equal to $Q$ and the adjoint potential $\overline{P}$ becomes equal to $Q_{x}.$ In that case, the equation in (2.8) reduces to
@@ 3 The direct scattering problem in the reflectionless case
where $\arg[k]$ denotes the argument function taking values in the interval $(-2\pi/3,4\pi/3).$ We use $\overline{\Omega_{1}},$ $\overline{\Omega_{2}},$ $\overline{\Omega_{3}},$ $\overline{\Omega_{4}}$ to denote the closures of the open sectors $\Omega_{1},$ $\Omega_{2},$ $\Omega_{3},$ $\Omega_{4},$ respectively, where we recall that the closures are obtained by adding the boundaries to the corresponding open sectors.
@@
$T_{\text{\rm{r}}}(k)=\displaystyle\frac{1}{T_{\text{\rm{l}}}(k)},\qquad k\in\mathbb{C}.$ (3.6)
It is known [3] that (2.8) has the particular solutions $m(k,x)$ and $n(k,x)$ with the respective $k$-domains $\overline{\Omega_{2}}$ and $\overline{\Omega_{4}},$ and in the reflectionless case they satisfy the spacial asymptotics given by
@@
Since $k$ appears as $k^{3}$ in (2.8), it follows that $\psi(zk,x)$ and $\psi(z^{2}k,x)$ are also solutions to (2.8) whenever $\psi(k,x)$ is a solution. The $k$-domain of $\psi(zk,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $2\pi/3$ radians around the origin of the complex $k$-plane. Similarly, the $k$-domain of $\psi(z^{2}k,x)$ is obtained from the $k$-domain of $\psi(k,x)$ by a clockwise rotation of $4\pi/3$ radians. This allows us to determine three linearly independent solutions to (2.8) at each $k$-value in the complex $k$-plane. Toward that goal, we divide $\Omega_{1}$ into the open subsectors $\Omega_{1}^{\text{\rm{up}}}$ and $\Omega_{1}^{\text{\rm{down}}}$ by using the negative real axis $\mathbb{R}^{-},$ and we divide the open sector $\Omega_{3}$ into the open subsectors $\Omega_{3}^{\text{\rm{down}}}$ and $\Omega_{3}^{\text{\rm{up}}}$ by using the positive real axis $\mathbb{R}^{+}.$ On the middle plot of Figure 3.1, we show the partitioning of the complex $k$-plane into the six open sectors $\Omega_{1}^{\text{\rm{up}}},$ $\Omega_{1}^{\text{\rm{down}}},$ $\Omega_{2},$ $\Omega_{3}^{\text{\rm{down}}},$ $\Omega_{3}^{\text{\rm{up}}},$ and $\Omega_{4}.$ We use -$\mathbb{R}^{-}$ and $\mathbb{R}^{+},$ respectively, to denote the directed half lines parametrized as
@@
For fixed real values of $x$ and $t,$ the large $k$-asymptotics of the basic solutions $f(k,x),$ $g(k,x),$ $m(k,x),$ and $n(k,x)$ are, respectively, given by [3]
@@
Using (3.11)–(3.14), we express the potentials $Q$ and $P$ in terms of $u_{1}(x)$ and $u_{2}(x)$ as
@@ 4 The bound states and the bound-state dependency constants
A bound state corresponds to a nontrivial solution to (2.8) which is square integrable in $x\in\mathbb{R}.$ If a bound state occurs at $k=k_{j}$ somewhere in the complex $k$-plane, then the number of linearly independent square-integrable solutions to (2.8) at $k=k_{j}$ determines the multiplicity of that bound state. In our paper, we only consider simple bound states, where the multiplicity of each bound state is equal to $1.$ We refer the reader to [4, 5, 6] for the treatment of bound states with multiplicities for various differential and difference equations. The analysis of solutions to (2.8) at $k=0$ is more challenging. When we analyze bound states for (2.8), we assume that a bound state does not occur at $k=0.$
We consider the bound states occurring at the zeros of $T_{\text{\rm{l}}}(k)^{-1}$ in $\Omega_{1}^{\text{\rm{down}}}$ or $\Omega_{1}^{\text{\rm{up}}}.$ Without loss of generality, it is enough to consider the bound states at the poles of $T_{\text{\rm{l}}}(k)$ in $\Omega_{1}^{\text{\rm{down}}}.$ It turns out [3] that the analysis of a bound state at $k=k_{j}$ with $k_{j}\in\Omega_{1}^{\text{\rm{down}}}$ yields useful information about the bound state at $k=k_{j}^{\ast}$ with $k_{j}^{\ast}\in\Omega_{1}^{\text{\rm{up}}}.$ We recall that we use an asterisk to denote complex conjugation.
@@
where we have the determinant of the relevant $3\times 3$ matrix on the right-hand side. The $3$-Wronskian of any three solutions to (2.8) at any particular $k$-value is zero if and only if those three solutions are linearly dependent. Furthermore, because of the absence of the term $\psi^{\prime\prime}$ in (2.8), the $3$-Wronskian of any three solutions is independent of $x$ and its value can be evaluated at any particular $x$-value. For example, using their asymptotics as $x\to\pm\infty,$ we evaluate the $3$-Wronskian of $f(k,x),$ $g(zk,x),$ and $n(z^{2}k,x)$ in $\overline{\Omega_{1}^{\text{\rm{down}}}}$ as
@@
Let us assume that the nonzero complex constant $k_{j}$ is located in the open sector $\Omega_{1}^{\text{\rm{down}}}$ and that it corresponds to a bound state. Then, we have $T_{\text{\rm{l}}}(k_{j})^{-1}=0.$ From (4.1) we see that the three solutions $f(k_{j},x),$ $g(zk_{j},x),$ and $n(z^{2}k_{j},x)$ to (2.8) are linearly dependent. This allows us [3] to express $f(k_{j},x)$ as a linear combination of $g(zk_{j},x)$ and $n(z^{2}k_{j},x)$ as
@@
for some complex-valued constants $D(k_{j})$ and $W(k_{j}).$
@@
where the nonzero complex constant $W(k_{j})$ corresponds to the dependency constant at the bound state with $k=k_{j}.$ On the other hand, again from the analysis in Section 3 of [3] it follows that $W(k_{j})=0$ if we have $\arg[k_{j}]\in[7\pi/6,4\pi/3).$ In that case, from (4.2) with $W(k_{j})=0$ we get
@@
with the nonzero complex constant $D(k_{j})$ corresponding to the dependency constant at the bound state with $k=k_{j}.$ If the potentials $Q$ and $P$ in (2.8) depend also on the parameter $t,$ then the solutions to (2.8) also depend on $t.$ Consequently, the dependency constants $W(k_{j})$ and $D(k_{j})$ appearing in (4.3) and (4.4) also depend on $t.$ For further analysis related to the bound states of (2.8), we refer the reader to [3].